# Multi-Source Spatial Knowledge Understanding for Immersive Visual Text-to-Speech

### v1

@@ Abstract
Visual Text-to-Speech (VTTS) aims to take the spatial environmental image as the prompt to synthesize the reverberation speech for the spoken content. Previous research focused on the RGB modality for global environmental modeling, overlooking the potential of multi-source spatial knowledge like depth, speaker position, and environmental semantics. To address the issues, we propose a novel multi-source spatial knowledge understanding scheme for immersive VTTS, termed MS2KU-VTTS. Specifically, we first prioritize RGB image as the dominant source and consider depth image, speaker position knowledge from object detection, and semantic captions from image understanding LLM as supplementary sources. Afterwards, we propose a serial interaction mechanism to deeply engage with both dominant and supplementary sources. The resulting multi-source knowledge is dynamically integrated based on their contributions. This enriched interaction and integration of multi-source spatial knowledge guides the speech generation model, enhancing the immersive spatial speech experience. Experimental results demonstrate that the MS2KU-VTTS surpasses existing baselines in generating immersive speech. Demos and code are available at: https://github.com/MS2KU-VTTS/MS2KU-VTTS.
@@
Visual Text-to-Speech (VTTS), Reverberation Speech, multi-source spatial knowledge, Acoustic Environment
Figure 1: The overview of MS2KU-VTTS, that includes the Multi-source Spatial Knowledge, the Dominant-Supplement Serial Interaction, the Dynamic Fusion and the Speech Generation.
@@ I Introduction
Visual Text-to-Speech (VTTS) aims to generate reverberation speech that corresponds to spoken content based on the spatial environmental image. As human-computer interaction advances, VTTS has become integral to intelligent systems, playing a crucial role in fields such as augmented reality and virtual reality.
@@
Unlike acoustic matching tasks that focus on adapting input speech to match a reference environment [12, 11, 7, 9], environmental Text-to-Speech (TTS) seeks to synthesize speech with the environmental characteristics of the reference based on given textual content [8]. For example, the VoiceLDM [14] utilize the pre-trained CLAP model [15] to map a textual or audio description into an environmental feature vector that controls the reverberation aspects of the generated audio. In addition, the Environment-aware TTS [8] design an environment embedding extractor that learns environmental features from the reference speech by reducing the distance between embeddings of the same environment while increasing the distance between embeddings from different environments. In a recent study, ViT-TTS [1], introduce a transformer-based visual-text encoder to extract global spatial knowledge from the RGB image. Building on these advancements, this paper focuses on employing visual information as the cue to generate reverberation audio for the targeted environment.
@@
However, prior research in VTTS predominantly focuses on the RGB modality as the singular source of spatial knowledge. Therefore, integrating multi-source spatial data—including depth imagery, speaker positioning, and spatial environmental semantics—could significantly enhance image comprehension. For example, the FEW-SHOTRIR [29] employs depth images to augment its grasp of environmental acoustics, markedly improving its capability to predict room impulse responses by embedding crucial spatial contexts and geometric details for precise audio simulations in unfamiliar settings. Moreover, the position of the speaker critically affects the perception of reverberation [16]. For instance, the ViGAS [17] utilizes speaker position data to refine the accuracy of spatial audio synthesis, aligning visual cues with auditory outputs, which substantially boosts the realism and spatial coherence of audio in novel-view acoustic synthesis tasks. Additionally, environment captions provide detailed spatial data on the layout and dimensions of settings, enhancing situational adaptability [14]. For instance, the AST-LDM [32] leverages textual descriptions of target environments as reference prompts to guide the acoustic scene transfer process, significantly expanding the model’s versatility and applicability in creating immersive audio experiences.
@@
To address above challenges, we propose a novel multi-source spatial knowledge understanding scheme for immersive VTTS, termed MS2KU-VTTS. Specifically, to expand the dimensions of environmental perception, we first prioritize the RGB image as the dominant source and consider depth image, speaker position knowledge from object detection as image-related supplement source, and semantic captions from image understanding LLM as semantic-related supplementary source. Afterwards, to deeply engage with both dominant and supplementary sources, we propose a Dominant-Supplement Serial Interaction mechanism to processes features from various data sources. It is through this sequential approach that not only is the model’s perception of spatial relationships deepened, but also, via complex interactions between dominant and supplementary sources, precise modeling of environmental reverberations is facilitated. In the end, to achieve a comprehensive understanding of the environment, the resulting multi-source knowledge is dynamically integrated based on their contributions. This enriched interaction and integration of multi-source spatial knowledge guides the speech generation model, enhancing the immersive spatial speech experience. We conducted comprehensive experiments on the effectiveness of MS2KU-VTTS. Both subjective and objective evaluations demonstrate that our model outperforms other state-of-the-art baseline systems significantly in terms of naturalness and perceptual quality.
@@ II-A Model Overview
As illustrated in the pipeline of Fig. 1, the proposed MS2KU-VTTS architecture comprises four components: 1) Multi-source Spatial Knowledge, constructed from both dominant and supplementary sources; 2) Dominant-Supplement Serial Interaction, which aggregates intra- and inter-modal information through RGB-Depth Interaction for comprehensive spatial understanding, including Speaker Position Enhanced Interaction, employing positional data to model environmental acoustics, and RGB-Semantic Interaction, acquiring spatial semantics from Gemini-generated captions; 3) Dynamic Fusion, wherein the multi-source knowledge is dynamically fused based on their contributions; and 4) Speech Generation, utilizing this integrated information to guide the synthesis of reverberant speech.
@@ II-B1 Dominant Source: RGB
II-B2 Image-related Supplement Source: Depth and Speaker Position
To capture the global knowledge within the Depth image, such as object arrangement, depth features are also extracted using a ResNet18, denoted as $F_{D}\in\mathcal{R}^{1\times D}$. To acquire fine-grained spatial knowledge, this work integrates the positional information of the speaker as additional guidance, inspired by previous studies [3, 13, 10, 17]. Specifically, we first detect all recognizable objects within the RGB image and to ascertain their corresponding anchor box coordinates. Subsequently, those coordinates categorized as “human”are selected and transformed into two-dimensional pixel coordinates $(x,y)$, which are normalized to [0, 1] by the image width and height, respectively. These coordinates are then mapped into a higher-dimensional space, the mathematical representation of which can be articulated as follows:
$\phi(x,y)=\Big{(}\sin(2^{k}\pi x),\cos(2^{k}\pi x),\sin(2^{k}\pi y),\cos(2^{k}\pi y)\Big{)}_{k=0}^{D-1},$ (1)
where $\phi(\cdot)$ denotes the high-dimensional embedding of the position of speaker, employing harmonic terms up to order $D$. The embeddings are subsequently processed by adaptive max pooling followed by a multi-layer perceptron to extract positional features, denoted as $F_{P}\in\mathcal{R}^{1\times D}$.
II-B3 Semantic-related Supplement Source: Spatial Image Caption
To efficiently fuse semantic information within the RGB image, Gemini [6] is employed to transform complex visual data into structured captions, which are subsequently processed by BERT [5] to extract spatial semantic features, denoted as $F_{S}$, where $F_{S}\in\mathcal{R}^{1\times D}$.
@@ II-C Dominant-Supplement Serial Interaction
As shown in the middle panel of Fig. 1, the Dominant-Supplement Serial Interaction (D-SSI) comprises three parts: 1) RGB-Depth Interaction, which aggregates information from dominant and image-related sources; 2) Speaker Position Enhanced Interaction, which models the relationship between source position and environmental acoustics; and 3) RGB-Semantic Interaction, which learns complex spatial semantics.
@@ II-C1 RGB-Depth Interaction
The RGB-Depth Interaction, inspired by LGPM [26], is introduced to achieve a more comprehensive representation of RGB and Depth features. It first refines the feature representations of each source through self-attention mechanisms, while simultaneously enabling cross-modal information fusion via cross-attention. After that, these representations are then aggregated to update the RGB and Depth features $F_{R}^{{}^{\prime}}$ and $F_{D}^{{}^{\prime}}$, which is formulated as:
$F_{D}^{{}^{\prime}}=F_{R}+\Phi_{sr}(F_{R},F_{R},F_{R})+\Phi_{cr}(F_{R},F_{D},F_{D}),$ (2)
$F_{R}^{{}^{\prime}}=F_{D}+\Phi_{sd}(F_{D},F_{D},F_{D})+\Phi_{cd}(F_{D},F_{R},F_{R}),$ (3)
where $\Phi_{s*}(\cdot)$ and $\Phi_{c*}(\cdot)$ denote the self-attention and cross-modal attention functions, respectively,
@@ II-C2 Speaker Position Enhanced Interaction
To gain a in-depth understanding of spatial knowledge within a visual scene, we propose an attention-based Speaker Position Enhanced Interaction. Specifically, the position-related visual features as $\mathcal{P}_{R/D}$ are computed, which can be formulated as:
$\mathcal{P}_{R/D}=F_{R/D}^{{}^{\prime}}\cdot\varphi((F_{P})^{T}\cdot F_{R/D}^{{}^{\prime}}),$ (4)
where $\varphi(\cdot)$ denotes the softmax function and $(\cdot)^{T}$ represents the transpose operator. Subsequently, the position-related and visual features are combined via concatenation:
$\mathcal{V}_{R/D}=\varphi(FC(Concat[F_{R/D}^{{}^{\prime}},\mathcal{P}_{R/D}])),$ (5)
@@ II-C3 RGB-Semantic Interaction
To capture spatial semantic from environment captions, we employ Spatial Semantic Attention by applying a multi-head attention mechanism:
$\mathcal{V}_{S}=\text{MultiHead}(F_{S},F_{R}^{{}^{\prime}},F_{R}^{{}^{\prime}}),$ (6)
where $V_{S}\in\mathbb{R}^{1\times D}$ is updated from $F_{S}$.
@@ II-D Dynamic Fusion
To effectively aggregate multi-source spatial knowledge, we employ a dynamic weighting approach inspired by MLA [24], which adjusts the contribution of each type of spatial knowledge based on its complexity, quantified by entropy. Higher entropy indicates greater difficulty in understanding, thereby reducing the relevance of that knowledge type. The entropy for each type as $u_{i}$ is computed as:
@@ III Experiments and Results
TABLE I: Comparison with baseline systems on the SoundSpaces-Speech dataset for Seen and Unseen scenarios. Includes both subjective (with 95% confidence intervals) and objective results across different systems. The MS2KU-VTTS significantly outperform the baselines with p-value $<0.001$.
@@ III-A Datasets
Utilizing the SoundSpaces-Speech dataset [16], we adhere to [12, 1] by removing out-of-view samples and partitioning the data into “test-unseen”—featuring room acoustics from novel scenes—and “test-seen”—comprising scenes encountered during training. The dataset encompasses 28,853 training, 1,441 validation, and 1,489 testing samples, each containing clean text, reverberant audio, and panoramic RGB-D images. Text sequences are converted into phonemes using an open-source tool [^1] https://github.com/Kyubyong/g2p. Consistent with [22, 20], spectrograms are extracted via FFT with parameters: size 1024, hop size 256, and window size 1024 samples; these are then transformed into 80-bin mel-spectrograms. Fundamental frequency (F0) is extracted from raw waveforms using Parselmouth [^2] https://github.com/YannickJadoul/Parselmouth.
@@ III-C Evaluation Metrics
We evaluate waveform quality using both objective metrics and subjective assessments, aligning with previous studies [21, 20]. For objective analysis, we randomly select 50 samples from the test set. Our evaluation metrics include: 1) Perceptual quality, assessed by human listeners and quantified by the Mean Opinion Score (MOS), which rates audio quality on 1 to 5 scale; 2) Room acoustics accuracy, measured by RT60 Error (RTE), which compares the reverberation time of predicted and target waveforms; 3) Mel Cepstral Distortion (MCD) [18], which measures the spectral distance between synthesized and reference mel-spectra.
@@ III-E Main Results
As presented in Table I, MS2KU-VTTS demonstrates superior performance, achieving optimal RTE (0.0745), MCD (4.5544), and MOS (3.875 $\pm$ 0.011) metrics in unseen environments. Statistical analysis indicates that MS2KU-VTTS significantly outperforms the baselines with a p-value less than 0.001. Although efficacy slightly decreases on the test-unseen set compared to seen scenarios—likely due to unfamiliar training conditions—the model still surpasses DiffSpeech [21] and ProDiff [20] across all metrics, particularly in RTE, underscoring the limitations of traditional TTS models with spatial environmental data. Our multi-source approach effectively captures comprehensive spatial knowledge, conferring significant advantages. The model surpasses VoiceLDM [14] due to our advanced Gemini-based image understanding, which outperforms VoiceLDM’s simpler environmental prompts. While ViT-TTS exceeds other baselines, it underperforms relative to our model in both environments, confirming the efficacy of our approach in generating environment-aligned reverberant speech.
III-F Ablation Results
TABLE II: Ablation study results. The terms Depth, Position, RGB, Semantics, R-DI, R-SI, and DF denote knowledge from Depth, Speaker Position, RGB, Semantics, RGB-Depth Interaction, Speaker Position Enhanced Interaction, and Dynamic Fusion, respectively. MS2KU-VTTS results are detailed in Table I.

### v2

@@ Abstract
Visual Text-to-Speech (VTTS) aims to take the environmental image as the prompt to synthesize reverberant speech for the spoken content. Previous works focus on the RGB modality for global environmental modeling, overlooking the potential of multi-source spatial knowledge like depth, speaker position, and environmental semantics. To address these issues, we propose a novel multi-source spatial knowledge understanding scheme for immersive VTTS, termed MS2KU-VTTS. Specifically, we first prioritize RGB image as the dominant source and consider depth image, speaker position knowledge from object detection, and Gemini-generated semantic captions as supplementary sources. Afterwards, we propose a serial interaction mechanism to effectively integrate both dominant and supplementary sources. The resulting multi-source knowledge is dynamically integrated based on the respective contributions of each source. This enriched interaction and integration of multi-source spatial knowledge guides the speech generation model, enhancing the immersive speech experience. Experimental results demonstrate that the MS2KU-VTTS surpasses existing baselines in generating immersive speech. Demos and code are available at: https://github.com/AI-S2-Lab/MS2KU-VTTS.
@@
Visual Text-to-Speech, Reverberant Speech, Multi-Source Spatial Knowledge
Figure 1: The overall architecture of MS2KU-VTTS.
@@ I Introduction
Visual Text-to-Speech (VTTS) aims to generate reverberant speech that matches spoken content based on the spatial environmental image. With advancements in human-computer interaction, VTTS has become an integral part of intelligent systems, playing a vital role in fields such as augmented reality and virtual reality.
@@
Unlike acoustic matching tasks, which focus on adapting input speech to match a reference environment [1, 2, 3, 4], VTTS aims to synthesize speech that incorporates the environmental characteristics of that environment based on given textual content [5]. For example, VoiceLDM [6] utilizes a pre-trained CLAP model [7] to transform text or audio descriptions into environmental feature vectors, which control the reverberation characteristics of the generated audio. In addition, Environment-aware TTS [8] designs an environment embedding extractor that learns environmental features from reference speech by optimizing embedding distances to distinguish between different environments. In a recent study, ViT-TTS [9] introduces a transformer-based visual-text encoder for extracting global spatial knowledge from an RGB image. Building on these advancements, this paper focuses on employing visual information as a guide to synthesize reverberant audio for the target environment.
@@
However, prior research in VTTS predominantly focuses on RGB images as the sole source of spatial knowledge. Therefore, integrating multi-source spatial data—including depth images, speaker position data, and spatial semantics—could significantly enhance spatial understanding. For example, FEW-SHOTRIR [10] employs depth images to enhance its understanding of environmental acoustics, significantly improving its ability to predict room impulse responses by incorporating key spatial contexts and geometric details for precise audio simulations in unfamiliar settings. Moreover, speaker position plays a crucial role in shaping the perception of reverberation [11]. For instance, ViGAS [12] utilizes speaker position data to refine the accuracy of spatial audio synthesis by aligning visual cues with auditory outputs. This approach enhances the realism and spatial coherence of audio in novel-view acoustic synthesis tasks. Additionally, environment captions provide detailed spatial knowledge about environmental layouts and dimensions, enhancing situational adaptability [6]. For instance, AST-LDM [13] leverages textual descriptions of target environments as reference prompts to guide the acoustic scene transfer process, enhancing the model’s versatility and applicability in generating immersive auditory experiences.
@@
To address the above challenges, we propose a novel multi-source spatial knowledge understanding scheme for immersive VTTS, termed MS2KU-VTTS. Specifically, to expand the dimensions of environmental perception, we first prioritize the RGB image as the dominant source and consider depth image, speaker position knowledge from object detection as the image-related supplement source, and semantic captions from Gemini Pro Vision [14] for image understanding as the semantic-related supplementary source.Afterwards, to deeply engage with both dominant and supplementary sources, we propose a Dominant-Supplement Serial Interaction mechanism to process features extracted from various knowledge sources. It is through this sequential approach that not only is the model’s perception of spatial relationships deepened, but also, via complex interactions between dominant and supplementary sources, precise modeling of environmental reverberations is facilitated. In the end, for comprehensive environmental understanding, the resulting multi-source knowledge is dynamically integrated based on each source’s contribution. This enriched interaction and integration of multi-source spatial knowledge guides the speech generation model, enhancing the immersive speech experience. We conducted extensive experiments to evaluate the effectiveness of MS2KU-VTTS. Both subjective and objective evaluations demonstrate that our model outperforms state-of-the-art baseline systems in terms of naturalness and perceptual quality.
@@ II-A Model Overview
As illustrated in the pipeline in Fig.1, the proposed MS2KU-VTTS architecture consists of four components: 1) Multi-source Spatial Knowledge, constructed from both dominant and supplementary sources; 2) Dominant-Supplement Serial Interaction aggregates intra- and inter-modal information through multiple mechanisms. Specifically, RGB-Depth Interaction enables comprehensive spatial understanding, Speaker Position Enhanced Interaction employs positional data to model environmental acoustics, and RGB-Semantic Interaction extracts spatial semantics from Gemini-generated captions; 3) Dynamic Fusion, wherein the multi-source knowledge is dynamically fused based on the contributions of each source; and 4) Speech Generation, which uses the integrated information to guide the synthesis of reverberant speech.
@@ II-B1 Dominant Source - RGB
II-B2 Image-related Supplement Source - Depth and Speaker Position
To capture the global knowledge within the depth image, such as object arrangement, depth features are also extracted using a ResNet18, denoted as $F_{D}\in\mathcal{R}^{1\times D}$. To acquire fine-grained spatial knowledge, this work integrates the positional information of the speaker as additional guidance, inspired by previous studies [16, 17, 18, 12]. Specifically, we first detect all recognizable objects within the RGB image to obtain their corresponding anchor box coordinates. Subsequently, those coordinates categorized as “human” are selected and transformed into two-dimensional pixel coordinates $(x,y)$, which are normalized to the range $[0,1]$ based on the RGB image width and height, respectively. These coordinates are then mapped into a higher-dimensional space. The mathematical representation is as follows:
$\begin{split}\phi(x,y)=\Big{(}&\sin(2^{k}\pi x),\cos(2^{k}\pi x),\\ &\sin(2^{k}\pi y),\cos(2^{k}\pi y)\Big{)}_{k=0}^{D-1}\end{split}$ (1)
where $\phi(\cdot)$ denotes the high-dimensional embedding of the position of speaker, employing harmonic terms up to a dimension of $D$. The embeddings are subsequently processed by adaptive max pooling followed by a multi-layer perceptron to extract positional features, denoted as $F_{P}\in\mathcal{R}^{1\times D}$.
II-B3 Semantic-related Supplement Source - Spatial Image Captions
To efficiently fuse semantic information within the RGB image, Gemini Pro Vision [14] is employed to transform complex visual information into structured captions, which are subsequently processed by BERT [19] to extract spatial semantic features, denoted as $F_{S}$, where $F_{S}\in\mathcal{R}^{1\times D}$.
@@ II-C Dominant-Supplement Serial Interaction
As shown in the middle panel of Fig.1, the Dominant-Supplement Serial Interaction (D-SSI) comprises three components: 1) RGB-Depth Interaction, which aggregates knowledge from dominant and image-related sources; 2) Speaker Position Enhanced Interaction, which models the relationship between source position and environmental acoustics; and 3) RGB-Semantic Interaction, which learns rich spatial semantics.
@@ II-C1 RGB-Depth Interaction
The RGB-Depth Interaction, inspired by LGPM [20], is introduced to achieve a more comprehensive representation of RGB and Depth features. It first refines the feature representations of each source through self-attention mechanisms, while simultaneously enabling cross-modal information fusion via cross-attention. After that, these representations are aggregated to update the RGB and Depth features $F_{R}^{\prime}$ and $F_{D}^{\prime}$, which are formulated as:
$F_{D}^{\prime}=F_{R}+\Phi_{sr}(F_{R},F_{R},F_{R})+\Phi_{cr}(F_{R},F_{D},F_{D}),$ (2)
$F_{R}^{\prime}=F_{D}+\Phi_{sd}(F_{D},F_{D},F_{D})+\Phi_{cd}(F_{D},F_{R},F_{R}),$ (3)
where $\Phi_{s*}(\cdot)$ and $\Phi_{c*}(\cdot)$ denote the self-attention and cross-modal attention functions, respectively.
@@ II-C2 Speaker Position Enhanced Interaction
To capture detailed spatial knowledge within a visual scene, we propose an attention-based Speaker Position Enhanced Interaction. Specifically, the position-related visual features, denoted as $\mathcal{P}_{R/D}$, computed as:
$\mathcal{P}_{R/D}=F_{R/D}^{\prime}\cdot\varphi((F_{P})^{T}\cdot F_{R/D}^{\prime}),$ (4)
where $\varphi(\cdot)$ denotes the softmax function and $(\cdot)^{T}$ represents the transpose operator. Subsequently, the position-related and visual features are combined through concatenation:
$\mathcal{V}_{R/D}=\varphi(FC(Concat[F_{R/D}^{\prime},\mathcal{P}_{R/D}])),$ (5)
@@ II-C3 RGB-Semantic Interaction
To capture spatial semantics from environmental captions, we employ Spatial Semantic Attention by applying a multi-head attention mechanism:
$\mathcal{V}_{S}=\text{MultiHead}(F_{S},F_{R}^{\prime},F_{R}^{\prime}),$ (6)
where $V_{S}\in\mathbb{R}^{1\times D}$ is derived by refining $F_{S}$.
@@ II-D Dynamic Fusion
To effectively aggregate multi-source spatial knowledge, we employ a dynamic weighting approach inspired by MLA [21], which adjusts the contribution of each type of spatial knowledge based on its complexity, quantified by entropy. The underlying assumption is that higher entropy indicates greater uncertainty or difficulty in understanding, which reduces the relative importance of that knowledge type in the fusion process. The entropy for each knowledge type, denoted as $u_{i}$, is computed as:
@@ III Experiments and Results
TABLE I: Comparison of MS2KU-VTTS and baselines on the SoundSpaces-Speech dataset for Seen and Unseen scenarios, including subjective (95% confidence intervals) and objective metrics. MS2KU-VTTS significantly outperforms baselines ($p<0.001$).
@@ III-A Datasets
Utilizing the SoundSpaces-Speech dataset [11], we follow the approach outlined in [1, 9], which involves removing out-of-view samples and partitioning the data into “test-unseen”—featuring room acoustics from novel scenes—and “test-seen”—comprising scenes encountered during training. The dataset encompasses 28,853 training, 1,441 validation, and 1,489 testing samples, each containing clean text, reverberant audio, and panoramic RGB-D images. Text sequences are converted into phonemes using an open-source tool [^1] https://github.com/Kyubyong/g2p. Consistent with [24, 22, 25, 26], spectrograms are extracted via FFT with parameters: size 1024, hop size 256, and window size 1024 samples; these are then transformed into 80-bin mel-spectrograms. Fundamental frequency is extracted from raw waveforms using Parselmouth [^2] https://github.com/YannickJadoul/Parselmouth.
@@ III-C Evaluation Metrics
We evaluate waveform quality using both objective metrics and subjective assessments, aligning with previous studies [23, 22]. For objective analysis, we randomly select 50 samples from the test set. Our evaluation metrics include: 1) Perceptual quality, assessed by human listeners and quantified by the Mean Opinion Score (MOS), which rates audio quality on a scale of 1 to 5; 2) Room acoustics accuracy, measured by RT60 Error (RTE), which compares the reverberation time of predicted and target waveforms; 3) Mel Cepstral Distortion (MCD) [28], which measures the spectral distance between synthesized and reference mel-spectra.
@@ III-E Main Results
As presented in Table I, MS2KU-VTTS demonstrates superior performance across all evaluation metrics, achieving the best RTE (0.0745), MCD (4.5544), and MOS (3.875 $\pm$ 0.011) in unseen environments. Statistical analysis indicates that MS2KU-VTTS significantly outperforms the baselines with a p-value less than 0.001. Although efficacy slightly decreases on the test-unseen set compared to seen scenarios—likely due to unfamiliar training conditions—the model still surpasses DiffSpeech [23] and ProDiff [22] across all metrics, particularly in RTE, underscoring the limitations of traditional TTS models with spatial environmental data. Our multi-source approach effectively captures comprehensive spatial knowledge, conferring significant advantages. The model surpasses VoiceLDM [6] due to our advanced Gemini-based image understanding, which outperforms VoiceLDM’s simpler environmental prompts. While ViT-TTS [9] exceeds other baselines, it underperforms relative to our model in both environments, confirming the efficacy of our approach in generating environment-aligned reverberant speech.
TABLE II: Ablation study results. Depth, Position, RGB, Semantics, R-DI, R-SI, and DF represent different knowledge components and interactions. Full MS2KU-VTTS results are in Table I.